Assume a monopolistic firm with a constant marginal cost (MC). The firm's demand schedule is given by table 1 below. Calculate the firm’s total revenue. Average revenue and marginal cost.

3 300 5 250

8 220

9 200

13 180 15 160 19 140 24 100 30 60

1.1 Find the firm's total revenue schedule, entering the data into the table where indicated. (10) 1.2 Use these data to determine

1.3 Assuming a constant marginal of R40, what output level and price will maximize the firm's profit? (5)

Question 2

Price stability is one of the main macroeconomic objectives hence monetary authorities pursue monetary policy to Suppose that you are a member of the Monetary Policy Committee of the SA Reserve Bank. The economy is experiencing a sharp increase in inflation.

2.1 What are the options for the SARB, what response do you recommend and why? [hint: tools of monetary policy] (15)

2.2 Despite its success, monetary policy has certain limitations and faces real-world complications.

 


1
Expert's answer
2021-05-17T18:19:26-0400

1.1 Find the rm's total revenue schedule, entering the data into the table where indicated.

"Total \space revenue = output \times price"



1.2 Use these data to determine the marginal revenue schedule



Marginal revenue when output is 5 = "\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{1250-900} { 5-3 }=\frac{350} { 2 }=175"

Marginal revenue when output is 8 = "\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{1760-1250} { 8-5 }=\frac{510} { 3 }=170"

Marginal revenue when output is 9 = "\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{1800-1760} { 9-8 }=\frac{40} { 1 }=40"

Marginal revenue when output is 13 = "\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{2340-1800} { 13-9 }=\frac{540} { 4 }=135"

Marginal revenue when output is 15 ="\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{2400-2340} { 15-3}=\frac{60} { 2 }=30"

Marginal revenue when output is 19 ="\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{2660-2400} { 19-15}=\frac{260} { 4 }=65"

Marginal revenue when output is 24 ="\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{2400-2660} { 24-19 }=\frac{-260} { 5 }=-52"

Marginal revenue when output is 30 ="\frac{Change \space in \space total \space revenue}{Change \space in \space output }=\frac{1800-2400} { 30-24 }=\frac{-600} { 6 }=-100"

1.3 Assuming a constant marginal of R40, what output level and price will maximize the firm's profit?

Prot maximizing output level is where marginal revenue equals marginal cost. Marginal revenue = Marginal cost = 40, at the quantity of output 9 and price 200R. Thus profit-maximizing output level is 9 units, and the profit maximizing price is 200R.



Question 2

Price stability is one of the main macroeconomic objectives; hence monetary authorities pursue a monetary policy to Suppose that you are a member of the Monetary Policy Committee of the SA Reserve Bank. The economy is experiencing a sharp increase in inflation.

2.1 What are the options for the SARB, what response do you recommend, and why? [hint: tools of monetary policy] (15)

Contractionary monetary policy is used when the economy is growing very rapidly or is facing increased risks of high inflation.

Central Bank decreases the money supply in contractionary monetary policy.



Central Bank decreases the money supply in the economy by the following methods:-


a) Increasing the discount rate - Discount rate is the rate at which Central Bank lends to the commercial banks. Increasing this rate will result in commercial banks borrowing less from the central bank, thus decreasing the money supply.


b) Increasing the reserve requirements - Each bank has to keep a certain percentage of its deposits as reserves; increasing reserve ratio will result in banks requiring to keep more reserves and thus lend less, thus decreasing the money supply.


c) Open market operations - Selling government bonds and securities to the open market (Commercial banks) and taking cash from the banks in exchange, thus decreasing the money supply.


A decrease in the money supply increases interest rates. At higher interest rates, firms are demotivated to borrow and thus borrows less and spend less.


Thus there is a decline in investment spending.

Investment spending is part of aggregate demand.

The decline in investment spending causes a decline in aggregate demand.

As Aggregate demand decreases, it causes the price levels in the economy to fall, thereby controlling inflation and eliminating the increased risk of high inflation.

2.2 Despite its success, monetary policy has certain limitations and faces real-world complications

Zero lower bound - When the economy faces a severe recession, the interest rates are already meager (near zero). The central bank cannot further lower the interest rates to motivate borrowing and spending in the economy. 


 


A liquidity trap occurs when the individuals and banks in the economy have no confidence in the economy. Interest rate cuts are not being passed on to consumers because banks are uncertain and afraid to lend. Thus interest rate cuts do not stimulate the economy, making expansionary monetary policy ineffective.


 


Interest rates and exchange rates - When the central bank increases the domestic interest rates. It also results in capital inflow, causing domestic currency appreciation, making exports expensive and less competitive, thus hurting domestic exporters, which the central bank never meant to do.


 


Time lags - When central banks change the base rates, it may take months for the new rates to be passed on to consumers by commercial banks; thus, monetary policy on the economy takes months after implementation.